OpenStudy (anonymous):

Add- The cubed root of x^6y^3 + The cubed root of x^3y^3

OpenStudy (hoblos):

= x^2y + xy = xy(x+1)

OpenStudy (anonymous):

Do mind telling me how u came to that.........

OpenStudy (hoblos):

\[\sqrt{x ^{3}} =x\] do u know this ?

OpenStudy (anonymous):

The product rule??

OpenStudy (hoblos):

i don't know its name but it is a rule and when u follow it u get the answer btw it is \[\sqrt[{3}]{x ^{3}} = x\] i typed it wrong before sorry

OpenStudy (anonymous):

The product rule??

OpenStudy (anonymous):

ok yeah thats the product rule and thanks that helped

OpenStudy (hoblos):

welcome :D